OpenStudy (anonymous):

Determine if the following expressions are equivalent? Justify your answer. 6x^2+7x-3 f(x)=_____________ and g(x)=3x(3x-1)-(1+9x^2) 2x+3

OpenStudy (anonymous):

i solved it numerically already, with a 2. it gives me the same answer, 5.

OpenStudy (helder_edwin):

\[ \large g(x)=3x(3x-1)-(1+9x^2)=9x^2-3x-1-9x^2=-(3x+1) \]

OpenStudy (anonymous):

i got that i just didnt know what to do with it

OpenStudy (helder_edwin):

using synthetic division for \(f(x)\): \[ \large \begin{array}{r|rrr} & 6 & 7 & -3\\ -3/2 & & -9 & 3\\ \hline & 6 & -2 & 0 \end{array} \] so \[ \large f(x)=\frac{(x+3/2)(6x-2)}{2x+3}= \frac{1/2(2x+3)(6x-2)}{2x+3}= \] \[ \large =\frac{1/2(6x-2)}{1}=3x-1 \]

OpenStudy (helder_edwin):

so \(f(x)=3x-1\) and \(g(x)=-3x-1\). \(f\equiv g\)??

OpenStudy (anonymous):

you confused me at the synthetic division part.

OpenStudy (anonymous):

thats not wat im learning right now.